Ryan has some bananas. The amount of bananas she has is two more than twice the amount of apples she has. If she has 29 bananas and apples total, how many of each does she have?
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39617)

Do you want to use one variable or two variables?

9 apples and 20 bananas

Answer by ikleyn(52787)

.

    Apples + bananas  = 29


      a    + (2a + 2) = 29


        3a + 2        = 29


           3a         = 29 - 2 = 27


            a                  = 27/3 = 9.


ANSWER.  9 apples and 29-9 = 20 bananas.

Solved.